# revealing an in-context neural scaling law

### v2

@@ A.3.2 Geometric Brownian motion
Geometric Brownian motion (GBM) (Oksendal, 2013) is a stochastic process that is commonly used in mathematical finance to model the trajectories of stock prices and other financial assets (Hull, 2021). A GBM is governed by the following SDE:
@@ A.4.1 Logistic map
The logistic map, first proposed as a discrete-time model for population growth, is one of the simplest dynamical systems that manifest chaotic behavior (Strogatz, 2015). It is governed by the following iterative equation:
@@
which may also be written using conditional distributions to reflect the deterministic nature of the system as $X_{t+1}|\{X_{t}=x_{t}\}\sim\delta_{f(x_{t})}$, where $\delta$ denotes the Dirac delta distribution. This conditional distribution is the ground truth transition function displayed in red in Figure 13. The parameter $r\in[1,4)$ controls the behavior of the system and is set to $r=3.9$. At this value, the dynamics are naturally confined within the interval $(0,1)$, and the system has no stable fixed points. Due to the chaotic nature of the system, two initial nearby trajectories diverge exponentially in time. This property allows us to sample multiple uncorrelated trajectories by using different initial conditions $x_{0}$, sampled uniformly in $(0,1)$.
@@ A.4.2 Lorenz system
The Lorenz system (Lorenz, 1963) is a three-dimensional (3D) dynamical system derived from a simplified model of convection rolls in the atmosphere. It consists of a system of three ordinary differential equations:
@@
When the $x$, $y$, and $z$ components are observed, the system is deterministic and Markovian; in the sense that a state vector $\vec{s}_{t}=(x_{t},y_{t},z_{t})$ at time $t$ fully determines the next state $\vec{s}_{t+1}$. However, if the $x$-component is the only one observed, then the system ceases to be Markovian but remains deterministic if one expands the state vector to include information from earlier states. Hence, Takens’ embedding theorem (Takens, 2006) guarantees that the observation of at most seven states of the series $x_{0:t}$ is sufficient to predict $x_{t+1}$. Finding the optimal number of states to reconstruct the system’s trajectory is an area of active research (Strogatz, 2015). Despite this apparent difficulty, LLaMA-13b can formulate increasingly accurate predictions of the series as it observes more states, as evidenced by the decaying loss curves plotted in Figure 15.
@@ A.5 Continuous State Space Visualization
One may naively remark upon the possibility that the in-context learning task for the Lorenz system and the logistic map could be rendered trivial if $x_{t+1}$ always falls close to $x_{t}$, in which case the LLM only needs to learn a static noisy distribution in order to decrease the loss. This is not the case with our experiments. In this section, we demonstrate the non-triviality of the learning tasks in Figures 17 and 17. In both cases, it is clear that the LLM has successfully learned to actively predict the expected mean position of the next state, and, in the logistic map example, the variance of the next state distribution as well. We note that the Lorenz system is simulated deterministically, hence the true next-state distribution is represented as a delta-function.
@@ A.6 In-context neural scaling law
Neural scaling laws (Kaplan et al., 2020) describe how the performance of trained neural networks, particularly language models, scales with changes in key factors such as model size ($N$), dataset size ($D$), and computational resources used for training (C). These laws are often observed as power-law relationships in the following form:
@@ A.7 Baselines for noisy logistic map and Markov chains
In this section, we compare LLM’s predictions against baseline models of known architectures, in order to understand the difficulty of the in-context learning task and make better sense of the Bhattacharyya loss. Specifically, we consider the following baseline models: unigram and bi-gram models for discrete Markov chains, and linear and non-linear autoregressive models with 1-step memory (AR1) for noisy logistic maps. The bi-gram model for the Markov chain has an unfair advantage since it is designed to model Markovian processes where the probability distribution of a token depends only on the previous token, ie., inferring the values of the transition matrix. The unigram model, on the other hand, models all tokens as drawn i.i.d. from the same distribution.
@@ A.8 Invariant measure and the early plateauing of in-context loss
While most datasets are well-described by the power laws, two loss curves — the Brownian motion and geometric Brownian motion — plateau early at a context length of about $10^{2}$, as shown in Figure 18. We attribute this early plateauing to the fact that the Brownian and geometric Brownian motions “wander out of distribution” at large time $t$, while all other dynamical systems studied in this paper converge to stable distributions (i.e., the invariant measure). A Markovian system (stochastic or deterministic) governed by a transition rule $P(x_{t+1}|x_{t})$ is said to have an invariant measure $\pi$ if
@@
If a system is initialized by $\pi(x)$ and evolves according to $P$, then the distribution of states at the next step will still follow $\pi(x)$. This property makes $\pi$ an invariant or stationary distribution for the system. It has been shown that the logistic map and Lorenz systems in the chaotic regime converge almost surely to their respective invariant measure, regardless of the initialization (Strogatz, 2015).
For discrete Markov chains governed by a transition matrix $p$, the stationary distribution is defined as a discrete probability mass function, denoted by $\vec{\pi}$, such that
@@
which is analogous to the continuous case described by Equation 12. By definition, any non-negative right eigenvector of $p$ with eigenvalue $\lambda=1$ is a stationary distribution of $p$. (Sethna, 2021) showed that a valid transition matrix has at least one stationary distribution. On the other hand, neither the Brownian nor the geometric Brownian motion has invariant distributions [^3] For stochastic systems, the invariant measure is sometimes referred to as the stationary distribution. on unbounded domains (e.g., when $\mathcal{X}=\mathbb{R}$). This can be seen from the marginalized distribution $P(x_{t})$ at time $t$. For the Brownian motion defined in Equation 8, the marginalized distribution of $x_{t}$ at time $t$ is a normal distribution:
@@
while for the geometric Brownian motion defined in Equation 9, the marginalized distribution of $x_{t}$ is a log-normal distribution (Crow & Shimizu, 1987):
@@
Both Equations 14 and 15 are time-dependent and do not converge to a stationary distribution in the limit $t\rightarrow\infty$. For the Brownian and geometric Brownian motions, the LLM might decide to only consider the most recent segment of time steps, and ignore the earlier data, which are in some sense “out of distribution”. This could explain the early plateauing of loss curves. Indeed, the classical neural scaling laws can be improved or broken if the scheduling of the training data shifts in distribution, as shown in (Sorscher et al., 2023). Different from (Sorscher et al., 2023; Lu et al., 2022), which alter the scheduling of data to achieve better learning curves that decrease faster with the size of training data, our experiments consider time series with pre-determined transition laws. We cannot tamper with the scheduling of our data to make it stationary without altering the underlying transition rules.
@@ A.9 Temperature and variance
The temperature $T$ is a hyper-parameter that controls the variance of the softmax output layer. Although most LLMs are trained at $T=1$, it is common practice to tune the temperature in the interval $T\in[0.8,1.2]$ during inference. Then, one can opt for increased diversity (high $T$), or better coherence (low $T$) in the generated output. The temperature hyper-parameter affects the uncertainty, or variance, in the Hierarchy-PDF extracted from the LLM. Figures 22 and 22 show how different temperatures change the shape of the Hierarchy-PDF. In both cases, higher temperature leads to higher variance in the PDF.
@@
We highlight the different refinement schemes used in these figures: for GBM, the PDF is refined to the last (third) digit near the mode, and left coarse elsewhere. This is because the true variance for GBM can span two orders of magnitude (see Figure 11), with most data points trapped in the low-variance region at small $X_{t}$. Hence, we require high precision to resolve these small variances in Figure 23. On the other hand, the noisy logistic map time series does not suffer from this issue, and thus we uniformly refine its PDF only up to the second digit.
@@
While the loss curves in our paper are calculated at $T=1$, the predicted $\sigma$ shown in Figures 11 and 6 are extracted at $T=0.7$. We performed a grid search on the temperature ranging from $T=0.3$ to $T=3$ (see Figures 25 and 25), and observed that $T=0.7$ consistently results in better prediction quality of the variance.

### v3

@@ A.3.2 Geometric Brownian motion
Geometric Brownian motion (GBM) Oksendal (2013) is a stochastic process that is commonly used in mathematical finance to model the trajectories of stock prices and other financial assets Hull (2021). A GBM is governed by the following SDE:
@@ A.4.1 Logistic map
The logistic map, first proposed as a discrete-time model for population growth, is one of the simplest dynamical systems that manifest chaotic behavior Strogatz (2015). It is governed by the following iterative equation:
@@
which may also be written using conditional distributions to reflect the deterministic nature of the system as $X_{t+1}|\{X_{t}=x_{t}\}\sim\delta_{f(x_{t})}$, where $\delta$ denotes the Dirac delta distribution. This conditional distribution is the ground truth transition function displayed in red in Figure 15. The parameter $r\in[1,4)$ controls the behavior of the system and is set to $r=3.9$. At this value, the dynamics are naturally confined within the interval $(0,1)$, and the system has no stable fixed points. Due to the chaotic nature of the system, two initial nearby trajectories diverge exponentially in time. This property allows us to sample multiple uncorrelated trajectories by using different initial conditions, $x_{0}$, sampled uniformly in $(0,1)$.
@@ A.4.2 Lorenz system
The Lorenz system Lorenz (1963) is a three-dimensional (3D) dynamical system derived from a simplified model of convection rolls in the atmosphere. It consists of a system of three ordinary differential equations:
@@
When the $x$, $y$, and $z$ components are observed, the system is deterministic and Markovian; in the sense that a state vector $\vec{s}_{t}=(x_{t},y_{t},z_{t})$ at time $t$ fully determines the next state $\vec{s}_{t+1}$. However, if the $x$-component is the only one observed, then the system ceases to be Markovian but remains deterministic if one expands the state vector to include information from earlier states. Hence, Takens’ embedding theorem Takens (2006) guarantees that the observation of at most seven states of the series $x_{0:t}$ is sufficient to predict $x_{t+1}$. Finding the optimal number of states to reconstruct the system’s trajectory is an area of active research Strogatz (2015). Despite this apparent difficulty, LLaMA-13b can formulate increasingly accurate predictions of the series as it observes more states, as evidenced by the decaying loss curves plotted in Figure 17.
@@ A.5 Continuous State Space Visualization
One may naively remark upon the possibility that the in-context learning task for the Lorenz system and the logistic map could be rendered trivial if $x_{t+1}$ always falls close to $x_{t}$, in which case the LLM only needs to learn a static noisy distribution in order to decrease the loss. This is not the case with our experiments. In this section, we demonstrate the non-triviality of the learning tasks in Figures 19 and 19. In both cases, it is clear that the LLM has successfully learned to actively predict the expected mean position of the next state, and, in the logistic map example, the variance of the next state distribution as well. We note that the Lorenz system is simulated deterministically, hence the true next-state distribution is represented as a delta-function.
@@ A.6 In-context neural scaling law
Neural scaling laws Kaplan et al. (2020) describe how the performance of trained neural networks, particularly language models, scales with changes in key factors such as model size ($N$), dataset size ($D$), and computational resources used for training (C). These laws are often observed as power-law relationships in the following form:
@@ A.7 Baselines for noisy logistic map and Markov chains
In this section, we compare LLM’s predictions against baseline models of known architectures, in order to understand the difficulty of the in-context learning task and make better sense of the Bhattacharyya loss. Specifically, we consider the following baseline models: unigram and bi-gram models for discrete Markov chains, and linear and non-linear autoregressive models with 1-step memory (AR1) for noisy logistic maps. The bi-gram model for the Markov chain has an unfair advantage since it is designed to model Markovian processes where the probability distribution of a token depends only on the previous token, i.e., inferring the values of the transition matrix. The unigram model, on the other hand, models all tokens as drawn i.i.d. from the same distribution.
@@ A.8 Invariant measure and the early plateauing of in-context loss
While most datasets are well-described by the power laws, two loss curves — the Brownian motion and geometric Brownian motion — plateau early at a context length of about $10^{2}$, as shown in Figure 20. We attribute this early plateauing to the fact that the Brownian and geometric Brownian motions “wander out of distribution" at large time $t$, while all other dynamical systems studied in this paper converge to stable distributions (i.e., the invariant measure). A Markovian system (stochastic or deterministic) governed by a transition rule $P(x_{t+1}|x_{t})$ is said to have an invariant measure $\pi$ if
@@
If a system is initialized by $\pi(x)$ and evolves according to $P$, then the distribution of states at the next step will still follow $\pi(x)$. This property makes $\pi$ an invariant or stationary distribution for the system. It has been shown that the logistic map and Lorenz systems in the chaotic regime converge almost surely to their respective invariant measures, regardless of the initialization Strogatz (2015).
For discrete Markov chains governed by a transition matrix, $p$, the stationary distribution is defined as a discrete probability mass function, denoted by $\vec{\pi}$, such that
@@
which is analogous to the continuous case described by Equation 12. By definition, any non-negative right eigenvector of $p$ with eigenvalue $\lambda=1$ is a stationary distribution of $p$. Sethna (2021) showed that a valid transition matrix has at least one stationary distribution. On the other hand, neither the Brownian nor the geometric Brownian motion has invariant distributions [^6] For stochastic systems, the invariant measure is sometimes referred to as the stationary distribution. on unbounded domains (e.g., when $\mathcal{X}=\mathbb{R}$). This can be seen from the marginalized distribution $P(x_{t})$ at time $t$. For the Brownian motion defined in Equation 8, the marginalized distribution of $x_{t}$ at time $t$ is a normal distribution:
@@
while for the geometric Brownian motion defined in Equation 9, the marginalized distribution of $x_{t}$ is a log-normal distribution Crow and Shimizu (1987):
@@
Both Equations 14 and 15 are time-dependent and do not converge to a stationary distribution in the limit $t\rightarrow\infty$. For the Brownian and geometric Brownian motions, the LLM might decide to only consider the most recent segment of time steps, and ignore the earlier data, which are in some sense “out of distribution”. This could explain the early plateauing of loss curves. Indeed, the classical neural scaling laws can be improved or broken if the scheduling of the training data shifts in distribution, as shown in Sorscher et al. (2023). Different from Sorscher et al. (2023); Lu et al. (2022), which alter the scheduling of data to achieve better learning curves that decrease faster with the size of training data, our experiments consider time series with pre-determined transition laws. Therefore, we cannot tamper with the scheduling of our data to make it stationary without altering the transition rules underlying the time series.
@@ A.9 Temperature and variance
The temperature $T$ is a hyper-parameter that controls the variance of the softmax output layer. Although most LLMs are trained at $T=1$, it is common practice to tune the temperature in the interval $T\in[0.8,1.2]$ during inference. Then, one can opt for increased diversity (high $T$), or better coherence (low $T$) in the generated output. The temperature hyper-parameter affects the uncertainty, or variance, in the Hierarchy-PDF extracted from the LLM. Figures 24 and 24 show how different temperatures change the shape of the Hierarchy-PDF. In both cases, higher temperature leads to higher variance in the PDF.
@@
We highlight the different refinement schemes used in these figures: for GBM, the PDF is refined to the last (third) digit near the mode, and left coarse elsewhere else. This is because the true variance for GBM can span two orders of magnitude (see Figure 13), with most data points trapped in the low-variance region at small $X_{t}$. Hence, we require high precision to resolve these small variances in Figure 25. On the other hand, the noisy logistic map time series does not suffer from this issue, and thus we uniformly refine its PDF only up to the second digit.
@@
While the loss curves in our paper are calculated at $T=1$, the predicted $\sigma$ shown in Figures 13 and 7 are extracted at $T=0.7$. We performed a grid search on the temperature ranging from $T=0.3$ to $T=3$ (see Figures 27 and 27), and observed that $T=0.7$ consistently results in better prediction quality of the variance.